OpenStudy (anonymous):

x^2+4y^2-16=0 what are the x and y intercepts

zepdrix (zepdrix):

Hley Gina :) The y-intercept is the location where we intercept the y axis. This happens when x=0.\[\Large\rm x^2+4y^2-16=0\]\[\Large\rm 0+4y^2-16=0\]What do you get for your y values?

OpenStudy (anonymous):

x^2+0-16=0?

zepdrix (zepdrix):

That's a good approach for finding your `x-intercepts`, by setting the y equal to 0. But let's them separately! :) Let's find our y-intercepts first. Need to solve for y here,\[\Large\rm 0+4y^2-16=0\]

zepdrix (zepdrix):

We would like to isolate our y. We'll start by doing the most simple steps first. We have a 16 being subtracted, so we'll do the opposite, we'll add 16 to each side, ya?

zepdrix (zepdrix):

\[\Large\rm 4y^2\cancel{-16+16}=0+16\]\[\Large\rm 4y^2=16\]

zepdrix (zepdrix):

What about the next step? We have a 4 multiplying our y^2. How do we undo multiplication? What's the opposite operation we can perform? :o

OpenStudy (anonymous):

4 divided by 16

zepdrix (zepdrix):

16 divided by 4* yes :)

OpenStudy (anonymous):

16/4 wich is y^2=4 then sqaure root both sides so y=2

zepdrix (zepdrix):

Well something fancy happens when we take the `square root` of a `square`. We end up with a plus or minus. We get two solutions! :) So we're actually going to get y=2, y=-2. k? :3

OpenStudy (anonymous):

right

zepdrix (zepdrix):

And then it looks like you already had the right approach for starting to find your x-intercepts:\[\Large\rm x^2+0-16=0\]

zepdrix (zepdrix):

Do some similar steps to what we did before, try to find some x values :D

OpenStudy (anonymous):

+16 to other side so it reads x^2=16

OpenStudy (anonymous):

square root each side to get +4 and -4

zepdrix (zepdrix):

Yayyy good job \c:/ Oh also, \(\Large\bf \color{#008353}{\text{Welcome to OpenStudy! :)}}\)